# Interpreting Key Mechanisms of Factual Recall in Transformer-Based Language Models

### v1

@@ Abstract
In this paper, we deeply explore the mechanisms employed by Transformer-based language models in factual recall tasks. In zero-shot scenarios, given a prompt like “The capital of France is,” task-specific attention heads extract the topic entity, such as “France,” from the context and pass it to subsequent MLPs to recall the required answer such as “Paris.” We introduce a novel analysis method aimed at decomposing the outputs of the MLP into components understandable by humans. Through this method, we quantify the function of the MLP layer following these task-specific heads. In the residual stream, it either erases or amplifies the information originating from individual heads. Moreover, it generates a component that redirects the residual stream towards the direction of its expected answer. These zero-shot mechanisms are also employed in few-shot scenarios. Additionally, we observed a widely existent anti-overconfidence mechanism in the final layer of models, which suppresses correct predictions. We mitigate this suppression by leveraging our interpretation to improve factual recall performance. Our interpretations have been evaluated across various language models, from the GPT-2 families to 1.3B OPT, and across tasks covering different domains of factual knowledge.
@@ 1 Introduction
Mechanistic interpretability research aims to reverse engineer deep neural networks into human-understandable algorithms or mechanisms (Wang et al., 2023; Olsson et al., 2022). This endeavor holds significant value for various applications, including providing insights into internal reasoning (Dutta et al., 2024; Kruthoff, 2024; Wendler et al., 2024), enriching knowledge (Meng et al., 2022; Hernandez et al., 2023), improving controllability (Yu et al., 2023; Chen et al., 2024), guiding future architecture and training paradigm designs (Elhage et al., 2022a; Lv et al., 2023; Reddy, 2024), and detecting safety issues (Shevlane et al., 2023). Despite the remarkable achievements made by language models (Brown et al., 2020; OpenAI, 2022, 2023; Touvron et al., 2023a, b) based on Transformer (Vaswani et al., 2017) in natural language understanding and generation, they remain mainly opaque to human, with many of their internal mechanisms understudied. In this paper, we delve into the mechanisms in factual recall tasks, which represent one of the cornerstone abilities of contemporary language models (Zhao et al., 2023; Zhang et al., 2023).
Recent observations made by Merullo et al. (2023) have delineated the behavior of language models when tasked with factual recall into two major phases as layers deepen: (1) Argument Formation: When presented with a query such as “The capital of France is”, the decoded top tokens from the residual stream gradually yield the name of the queried country, in this case, “France.” The authors likened this process to the model forming an implicit function akin to “get_capital(x),” where “France” is the argument for the implicit function. (2) Function Application: As layers deepen, the top tokens decoded from the residual stream transition from the country name to the capital name, such as shifting from “France” to “Paris.” The authors described this transition as the application of the implicit function and highlighted the crucial role of MLPs in this phase.
Although these observations lay a valuable groundwork, several unresolved research questions remain, which are essential for achieving a deep understanding of the factual recall mechanism in language models. For instance, how does the model extract the argument from the context and then pass it to the so-called “function?” What exactly constitutes “function application,” and how does it correlate with MLPs? Moreover, (Merullo et al., 2023) primarily focused on a one-shot setting. What about the factual recall mechanism of language models in basic zero-shot settings or when provided with more shots?
@@
In this paper, we delve into the mechanistic interpretability of these crucial questions and offer several innovative insights. We investigated the scalability of Transformer-based language models, ranging from GPT-2 small, medium, and large (Radford et al., 2019), to OPT-1.3B (Zhang et al., 2022), across various factual recall tasks encompassing diverse domains of knowledge. Based on our analysis, we outline the sequential steps by which a language model achieves factual recall in zero-shot [^1] For readers unfamiliar with mechanistic interpretability research, we suggest referring to Section 2 for necessary background information beforehand.:
(1) First, task semantics formed in shallow layers activate some task-specific attention heads in the mid-to-deep layers. These attention heads possess a QK matrix inherently sensitive to tokens related to specific topics, such as the name of a country. They attend to these topic tokens and move them to the final position of the residual stream. This mechanism enables the model to extract the “argument” from the context and pass it for further processing, e.g., the function application. Additionally, we observed that some attention heads’ OV matrix can directly map the argument to the desired outputs without further processing via MLPs. This mapping can be viewed as a partial function application.
(2) Given that all attention heads’ outputs are aggregated equally before being added to the residual stream, the subsequent MLP serves as an “activation” for each head’s output. It can erase or amplify the outputs of individual heads by generating vectors that either align with or oppose the direction of heads’ outputs. Through this process, the passed argument by task-specific heads “stands out” within the residual stream.
(3) In addition to activating attention heads, the MLP’s output also incorporates a task-aware component that directs the residual stream towards the direction of the target token’s unembedding vector. We demonstrated that the component’s addition into the residual stream accomplishes the “function application.”
We illustrate these three steps in Figure 1. Furthermore, we observed that various models suppresses correct predictions in the final layer. This is achieved by integrating frequent tokens into the residual stream via attention heads and employing the MLP to steer the residual stream towards a “mean” token over the training corpus. This phenomenon appears to be a universal anti-overconfidence mechanism adopted by various models to evade a significant training loss, and it persists irrespective of the task, model, and the number of in-context demonstrations.
@@
1. In zero-shot scenarios, we offer a detailed interpretation of the “argument passing” and “function application” mechanisms employed in Transformer-based language models for factual recall tasks. The mechanisms we detected still work in few-shot scenarios.
@@
3. We observed a universal anti-overconfidence mechanism in the final layer of models. Building on our interpretation, we devise strategies to mitigate this anti-overconfidence mechanism, enhancing factual recall performance.
@@ 2.1 A Residual-Centric Perspective of the Transformer
Elhage et al. (2021) presented a mathematical framework for comprehending the Transformer architecture, which we follow in this paper. In Transformer models, the residual connections (He et al., 2016) serve as the data bus for information flow. Each Transformer layer comprises multiple attention heads and an MLP layer, each of which reads from and writes to this data bus.
@@
Figure 1 illustrates the Transformer layer $l$ from this perspective. Elhage et al. (2021) reformulated attention heads for a better theoretical understanding of models, which is equivalent to that presented in (Vaswani et al., 2017). This reformulation can be represented as follows:
@@
2.2 Circuit Discovery
Discovering Circuit by Ablation
Figure 2: Subfigures (a), (b), and (c) showcase a toy diagram of causal mediation analysis in discovering important circuits in a neural network. Colors distinguish different values of the node. This example is inspired by the case described in §2.3, (Lieberum et al., 2023). Subfigure (d) illustrates an activation patching example in studying the node affecting the correct capital city prediction, as detailed in §2.2.
Let us conceptualize the neural network as a causal graph, where each node represents either neurons (Elhage et al., 2022a, b; Makelov et al., 2024) or module activations (Elhage et al., 2021). When there is a direct connection between two nodes in the network, they are linked by an edge in the graph. The graph is considered causal because intervening in one node’s output can causally influence the subsequent nodes that rely on its output as input. A circuit is a subgraph of the neural network (Olah et al., 2020). Research indicates that only a few nodes are working given a particular task Wang et al. (2023); Merullo et al. (2024). The objective of circuit discovery is to uncover the circuit comprising these activated nodes.
The prevailing approach for circuit discovery is grounded in causal mediation analysis (Pearl, 2001; Vig et al., 2020). Figure 2 illustrates a toy example to elucidate this process. Suppose we want to discern the impact of node $g$ on node $f$. When we intervene in $g$’s output (e.g., by knocking out or corrupting it), $f$ is affected by both the change from $g$ and $g$’s indirect intervention via $z$, as shown in Figure 2(b). In a more nuanced approach, we could maintain $z$’s outputs consistent to measure $g$’s “direct effect” to $f$, as shown in Figure 2(c). By conducting such ablation studies, we can first identify nodes that significant affect the neural network’s final outputs. Then, we can pinpoint which nodes influence these influential nodes through iterative backward tracing. This iterative process forms the fundamental concept behind many circuit discovery methods (Wang et al., 2023; Lieberum et al., 2023; Conmy et al., 2023).
@@ 2.2 Circuit Discovery
“Activation patching” is a widely employed technique for implementing the interventions in module activations (Wang et al., 2023; Meng et al., 2022; Vig et al., 2020). Suppose we have an original prompt, denoted as $p_{\text{original}}$, such as “The capital of $X$ is,” with the expected answer being $Y$. The activations of node $f$ in the neural network given this prompt is represented as $f(p_{\text{original}})$. Suppose we aim to pinpoint the node responsible for extracting country identity, which consequently influences the final node $f$, i.e., $\textbf{r}_{\text{post},t=-1}$ in the final layer. We might intervene on $g$, a preceding node of $f$ in the causal graph, at the country name position [^2] Patching the country name serves as a mere example. If the study aims to identify a node responsible for task information, the “capital” should be patched instead. Let us say we have $N$ intervention prompts sharing the template with $p_{\text{original}}$ but differing in keywords. For instance, each intervention prompt $p_{\text{intervention},n}$ takes the form “The capital of $X_{n}$ is,” where $X_{n}$ is sampled from a set of country names distinct from $X$. By substituting $g(p_{\text{original}})$ with $\frac{1}{N}\sum^{N}_{n=1}g(p_{\text{intervention},n})$, we corrupt the country identity in $g(p_{\text{original}})$. This process can be formalized as:
$\hat{g}(p_{\text{original}})=g(p_{\text{original}}\ |\ \texttt{do}(g(p_{\text{original}})=\frac{1}{N}\sum^{N}_{n=1}g(p_{\text{intervention},n}))).$ (3)
@@
Here, the operation $\texttt{do}(\cdot)$, a standard notation in causality (Pearl, 2009), indicates the replacement of one value with another. When the node $f$ takes $\hat{g}(p_{\text{original}})$ as inputs, we denote its current value as $\hat{f}$. If certain metrics indicating the probability of $Y$ decoded from $\hat{f}$ are lower compared to those from $f$, we can deduce that the edge from $g$ to $f$ has a positive impact on correct predictions, and vice versa [^3] The patching method introduced here is called “mean-ablate.” While there are some variants like “zero-ablate” or “Gaussian noising,” recent evidence (Wang et al., 2023; Zhang and Nanda, 2024) suggests that they may not be optimal. Consequently, we do not delve into them in detail.. The patching procedure is illustrated in Figure 2(d).
The logit lens technique (nostalgebraist, 2020) is widely used to evaluate the effects of patching the path from $g$ to $f$. It operates by decoding $f$ to obtain the logits corresponding to the correct answer $Y$. To implement this, we introduce a function called “Unembed,” defined as follows:
$\texttt{Unembed}(\cdot)=\textbf{W}_{U}\texttt{LN}(\cdot).$ (4)
Here, $\texttt{LN}(\cdot)$ represents the normalization in the final layer. $\textbf{W}_{U}\in\mathbb{R}^{V\times d_{\text{model}}}$, where $V$ is the vocabulary size, denotes the unembedding matrix (i.e., the weight of the language model head). We obtain $\pi(Y,f)$, the logit value of $Y$ decoded from $f$, as follows:
$\pi(Y,f)_{t=-1}=\left(\texttt{Unembed}(f)[Y]-\frac{1}{V}\sum^{V}_{v}\texttt{Unembed}(f)[v]\right)_{t=-1}.$ (5)
Here, $V$ represents the vocabulary size, and [$Y$] follows the Python syntax, selecting the logit of the token $Y$. We center the logits in order to measure the net effect of patching. The subscript $t=-1$ indicates our focus solely on the logit change at the last position where the current prediction occurs. Now, we can measure the impact of patching by evaluating the degree of change in logits:
$\Delta\pi(Y,f)_{t=-1}=(\frac{\pi(Y,\hat{f})_{t=-1}}{\pi(Y,f)_{t=-1}}-1)\cdot 100\%.$ (6)
@@ 3 How Does the Language Model Do Factual Recall Tasks in Zero-Shot
Recall that Merullo et al. (2023) observed that during factual recall, the language model first forms an “argument”. Subsequently, it passes the argument to an implicit function for deriving the final answer. In this section, we explore the following unresolved research questions within a zero-shot scenario:
@@ 1. How does the model discern the “argument” from context and pass it to the “implicit function?”
2. What exactly is the “function application”, and how does it relate to MLPs?
@@ 3.1 Experiment Setup
We study Transformer-based language models, ranging in scale from the 117M GPT-2 small (Radford et al., 2019), GPT-2 medium and large, up to the OPT-1.3B (Zhang et al., 2022). Our study involved two factual recall tasks concerning country-capital and product-developer associations, requiring different domain knowledge. The main text details our experiments conducted with GPT-2 small, which has 12 layers and 12 heads in each layer, using the country-capital task. Other experiments are presented in the appendix. Unless otherwise specified, the interpretations below are generally validated across models and tasks.
For the country-capital task, we prompt the language model, asking it to recall the capital of a given country. For instance, a prompt would be framed as “As we all know, the capital of $X$ is.” Here, $X$ denotes the country in question, and we denote the corresponding capital of $X$ by $Y$. To ensure the robustness of our experiments and to mitigate any bias stemming from specific prompt templates, we devise a set of 15 distinct prompts. We construct a set containing 15 country-capital pairs. To generate a test dataset, we fill each template with various country names, resulting in a dataset comprising $N=225$ samples. Data details can be found in appendix A.

### v4

@@ Abstract
In this paper, we delve into several mechanisms employed by Transformer-based language models (LLMs) for factual recall tasks. We outline a pipeline consisting of three major steps: (1) Given a prompt “The capital of France is,” task-specific attention heads extract the topic token, such as “France,” from the context and pass it to subsequent MLPs. (2) As attention heads’ outputs are aggregated with equal weight and added to the residual stream, the subsequent MLP acts as an “activation,” which either erases or amplifies the information originating from individual heads. As a result, the topic token “France” stands out in the residual stream. (3) A deep MLP takes “France” and generates a component that redirects the residual stream towards the direction of the correct answer, i.e., “Paris.” This procedure is akin to applying an implicit function such as “get_capital($X$),” and the argument $X$ is the topic token information passed by attention heads. To achieve the above quantitative and qualitative analysis for MLPs, we proposed a novel analytic method aimed at decomposing the outputs of the MLP into components understandable by humans. Additionally, we observed a universal anti-overconfidence mechanism in the final layer of models, which suppresses correct predictions. We mitigate this suppression by leveraging our interpretation to improve factual recall confidence. The above interpretations are evaluated across diverse tasks spanning various domains of factual knowledge, using various language models from the GPT-2 families, 1.3B OPT, up to 7B Llama-2, and in both zero- and few-shot setups.
@@ 1 Introduction
Mechanistic interpretability research aims to reverse engineer deep neural networks into human-understandable algorithms or mechanisms [39, 25]. This endeavor holds significant value for various applications, including providing insights into internal reasoning [5, 15, 41], enriching knowledge [20, 13], improving controllability [43, 2], guiding future architecture and training paradigm designs [6, 32], and detecting safety issues [33]. Despite the remarkable achievements made by language models [1, 26, 27, 34, 35] based on Transformer [37] in natural language understanding and generation, they remain mainly opaque to human, with many of their internal mechanisms understudied. In this paper, we delve into the mechanisms in factual recall tasks, which represent one of the cornerstone abilities of contemporary language models [47, 46].
Merullo et al. [21] delineated the behavior of language models in factual recall tasks into two major phases as layers deepen: (1) Argument Formation: When presented with a query such as “The capital of France is,” the decoded top-probable tokens from the residual stream gradually yield the name of the queried country, in this case, “France.” The authors likened this process to the model forming the argument $X$ for an implicit function such as “get_capital($X$).” (2) Function Application: As layers deepen, the top-probable tokens decoded from the residual stream transition from the country name to the capital name, such as from “France” to “Paris.” The authors described this transition as the application of the implicit function and assumed the crucial role of MLPs in this phase.
Although these observations lay a valuable groundwork, several unresolved research questions remain, which are essential for achieving a deep understanding of the factual recall mechanisms in language models. For instance, how does the model extract the argument from the context and then pass it to the so-called “function?” What exactly constitutes “function application,” and how does it correlate with MLPs? Moreover, Merullo et al. [21] primarily focused on a one-shot setting. What about the mechanisms of language models in basic zero-shot settings or when provided with more shots?
@@
In this paper, we investigate the mechanistic interpretability of these crucial questions and offer several innovative insights. For readers unfamiliar with mechanistic interpretability research, we suggest referring to Section 2 for necessary backgrounds. We investigated Transformer-based language models scaling from GPT-2 small, medium, and large [31], 1.3B OPT [45], to 7B Llama-2 [35], across various factual recall tasks encompassing diverse domains of knowledge. Based on our analysis, we outline the following steps by which a language model achieves factual recall.
(1) Some task-specific attention heads in the mid-to-deep layers possess weights inherently sensitive to topic tokens, such as the name of a country, which move these tokens to the final position of the residual stream. This mechanism enables the model to extract the “argument” from the context and pass it to the later “function application.” Additionally, some attention heads directly map the argument to the desired outputs, which can be viewed as a partial function application.
(2) The subsequent MLP serves as an “activation” to make the output of task-specific heads stand out in the residual stream, as the output of heads that pass arguments is aggregated with other heads with equal weight. This MLP can erase or amplify the outputs of individual heads by generating vectors that either align with or oppose the direction of heads’ outputs.
(3) Some deep MLP’s output incorporates a task-aware component that directs the residual stream towards the direction of the target token’s unembedding vector. The component’s addition into the residual stream accomplishes the “function application.” The above insights regarding MLPs come from our novel analytical approach, which deconstructs the MLP’s output into a linear combination of attention heads’ outputs as well as an additional component that encodes the MLP’s functionality.
We illustrate these three steps in Figure 1. These mechanisms consistently work in zero-, one-, and few-shot scenarios. Beyond answering these questions, we have observed that language models universally suppress correct predictions in the final layer. This is achieved by integrating frequent tokens into the residual stream via attention heads and employing the MLP to steer the residual stream towards an “average” token over the training corpus. This phenomenon appears to be a universal anti-overconfidence mechanism adopted by various models to evade significant training loss and persists irrespective of the task, model, and the number of in-context demonstrations.
@@
1. We examine various models, tasks, and ICL setups, offer a detailed and universal interpretation of the “argument passing” and “function application” mechanisms employed in Transformer-based language models for factual recall tasks.
@@
3. We observed a universal anti-overconfidence mechanism in the final layer of models. Building on our interpretation, we devise strategies to mitigate this anti-overconfidence mechanism, enhancing factual recall confidence.
@@ 2.1 A residual-centric perspective of the Transformer
Figure 1 illustrates the Transformer layer $l$ from this perspective. Elhage et al. [8] reformulated attention heads for a better theoretical understanding of models, which is equivalent to that presented in [37]. This reformulation can be represented as follows:
@@
2.2 Discovery of influential modules
Causal mediation analysis
Research indicates that only a few modules (e.g., individual heads, an MLP, etc.) work for a particular task [39, 22]. Circuit discovery research aims to uncover these activated modules, grounded in causal mediation analysis [29, 38].
Let us conceptualize the neural network as a causal graph, where each node represents either neurons [6, 7, 18] or module activations [8]. When there is a direct connection between two nodes in the network, they are linked by an edge in the graph. The graph is considered causal because intervening in one node’s output can causally influence the subsequent nodes that rely on its output as input. A neural network subgraph is termed a circuit [24]. In Figure 2, a toy example elucidates this process. Suppose we want to discern the impact of node $g$ on node $f$. When we intervene in $g$’s output (e.g., by knocking out or corrupting it), $f$ is affected by both the change from $g$ and $g$’s indirect intervention via $z$, as shown in Figure 2(b). More precisely, we could maintain $z$’s outputs consistent to measure $g$’s “direct effect” to $f$, as shown in Figure 2(c). By conducting such ablation studies, we can first identify nodes that significant affect the neural network’s final outputs.
@@ 2.2 Discovery of influential modules
Figure 2: Subfigures (a), (b), and (c) are toy diagrams of causal mediation analysis in discovering important circuits in a neural network. Colors distinguish different node values. Subfigure (d) illustrates an activation patching example in studying the node affecting the correct capital city prediction, as detailed in §2.2.
@@
“Activation patching” is a widely employed technique for implementing the interventions in module activations [39, 20, 38]. Suppose we have an original prompt, denoted as $p_{\text{original}}$, such as “The capital of $X$ is,” with the expected answer being $Y$. The activations of node $f$ in the neural network given this prompt is represented as $f(p_{\text{original}})$. Suppose we aim to uncover the node responsible for extracting country identity, which consequently influences the final node $f$, i.e., $\textbf{r}_{\text{post},t=-1}$ in the final layer. A feasible choice is to intervene on $g$, a preceding node of $f$ in the causal graph, at the country name position.
Let us say we have $N$ intervention prompts sharing the template with $p_{\text{original}}$ but differing in keywords. For instance, each intervention prompt $p_{\text{intervention},n}$ takes the form “The capital of $X_{n}$ is,” where $X_{n}$ is sampled from a set of country names distinct from $X$. By substituting $g(p_{\text{original}})$ with intervention information, we corrupt the country identity in $g(p_{\text{original}})$. This process can be formalized as:
$\hat{g}(p_{\text{original}})=g\left(p_{\text{original}}\ |\ \texttt{do}(g(p_{\text{original}})=\frac{1}{N}\sum^{N}_{n=1}g(p_{\text{intervention},n}))\right).$ (3)
@@
The operation $\texttt{do}(\cdot)$, a standard notation in causality [30], indicates the replacement of one value with another. When the node $f$ takes $\hat{g}(p_{\text{original}})$ as inputs, we denote its current value as $\hat{f}$. If certain metrics indicating the probability of $Y$ decoded from $\hat{f}$ are lower than those from $f$, we can deduce that the edge from $g$ to $f$ positively impacts correct predictions, and vice versa. The above patching procedure is illustrated in Figure 2(d).
A feasible metric is the difference in logit values, $\pi(Y,f)$, for our target token $Y$, decoded from node $f$ before and after patching node $g$. This metric, denoted $\Delta\pi(Y,f)$, can be formalized as follows:
$\quad\quad\quad\Delta\pi(Y,f)_{t}=(\frac{\pi(Y,\hat{f})_{t}}{\pi(Y,f)_{t}}-1)\cdot 100\%,$ (4)
where $\pi(Y,f)_{t}$ is defined as:
$\pi(Y,f)_{t}=\left(\texttt{Unembed}_{t}(f)[Y]-\frac{1}{V}\sum^{V}_{v}\texttt{Unembed}_{t}(f)[v]\right),$ (5)
and $\texttt{Unembed}(\cdot)$ is computed using $\textbf{W}_{U}\texttt{LN}(\cdot)$. Here, $\texttt{LN}(\cdot)$ represents the normalization in the final layer. $\textbf{W}_{U}\in\mathbb{R}^{V\times d_{\text{model}}}$, where $V$ is the vocabulary size, denotes the unembedding matrix (i.e., the weight of the language model head). The notation [$Y$] follows Python syntax, indicating the selection of $Y$-th item. The subscript $t$ indicates the focus position, set to $-1$ by default in this paper, representing the last position where the current prediction occurs. Without ambiguity, this subscript is omitted. In Eq. 5, we center the logits to measure the net effect of path patching.
@@ 3 How does the language model do factual recall tasks in zero-shot
Recall that in [21], authors observed that during factual recall, the language model first forms an “argument.” Subsequently, it passes the argument to an implicit function for deriving the final answer. In this section, we explore the following unresolved research questions within a zero-shot scenario:
@@ 1. How does the model discern the “argument” from context and pass it to the “implicit function?”
2. What exactly is the “function application,” and how does it relate to MLPs?
@@ 3.1 Experiment setup
We study Transformer-based language models, scaling from the 117M GPT-2 small [31], GPT-2 medium and large, OPT-1.3B [45], to the Llama-2 7B [35]. Our study involved two factual recall tasks concerning country-capital and product-developer associations, requiring different domain knowledge. The main text details our experiments on GPT-2 small, which has 12 layers and 12 heads in each layer, using the country-capital task. Other experiments are presented in appendices. Unless specified, the interpretations below are generally validated across models and tasks.
For the country-capital task, we prompt the language model, asking it to recall the capital of a given country. For instance, a prompt would be framed as “As we all know, the capital of $X$ is.” Here, $X$ denotes the country in question, and we denote the corresponding capital name of $X$ by $Y$. To ensure the robustness of our experiments and to mitigate any bias stemming from specific prompt templates, we devise a set of 15 distinct prompts. We construct a set containing 15 country-capital pairs. To generate a test dataset, we fill each template with various country names, resulting in a dataset comprising $N=225$ samples. Data details can be found in Appendix C.
Figure 3: (a) Probabilities of $X$ and $Y$ decoded at each layer. The shaded regions indicate variances. (b) The effect of patching $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post},t=-1}$. (c) Value weighted attention pattern of L9H8 and L10H0. (d) L9H8 and L10H0’s attention to $X$ is proportional to the projection value of their output along $\textbf{W}_{U}[X]$, indicating they are moving country names to the final position.